### v1

# Your Paper Title Here

Abstract
Machine learning models trained on Earth observation data, such as satellite imagery, have demonstrated significant promise in predicting household-level wealth indices, enabling the creation of high-resolution wealth maps that can be leveraged across multiple causal trials while addressing chronic data scarcity in global development research. However, because standard training objectives prioritize overall predictive accuracy, these predictions inherently suffer from shrinkage toward the mean, leading to attenuated estimates of causal treatment effects and limiting their utility in policy evaluations. Existing debiasing methods, such as Prediction-Powered Inference (PPI), can handle this attenuation bias but require additional fresh ground-truth data at the downstream stage of causal inference, which restricts their applicability in data-scarce environments. In this paper, we introduce and evaluate two correction methods—linear calibration correction and Tweedie’s correction—that substantially reduce prediction bias without relying on newly collected labeled data. Linear calibration (LCC) corrects bias through a straightforward linear transformation derived from held-out calibration data, whereas Tweedie’s correction leverages empirical Bayes principles to directly address shrinkage-induced biases by exploiting score functions derived from evaluating the model’s learning patterns. Through analytical exercises and experiments using Demographic and Health Survey (DHS) data, we demonstrate that both proposed methods meet or outperform existing approaches that either require (a) adjustments to training pipelines or (b) additional labeled data, achieving significant reductions in attenuation bias in data-scarce environments. These approaches may represent a promising avenue for improving the reliability of causal inference when direct outcome measures are limited or unavailable, enabling a “one map, many trials” paradigm where a single upstream data creation team produces predictions usable by many downstream teams across diverse ML pipelines.
Introduction
Machine learning (ML) models trained on Earth observation (EO) data have recently demonstrated impressive performance in predicting household-level wealth indices, such as the International Wealth Index (IWI), with reported $R^{2}$ values approaching 0.80 (Pettersson et al., 2023; Jean et al., 2016; Daoud et al., 2023; Chi et al., 2022). These EO-ML models offer a much-needed solution to longstanding data scarcity issues in global development, enabling researchers to generate high-resolution, retrospective “wealth maps” in data-poor regions. But while these predictions are promising in the EO-ML upstream task, a key question remains (Jerzak et al., 2023a, b): the value of EO-ML-derived data hinges on whether domain researchers, in the downstream phase, can use these data to estimate causal effects for their specific application area downstream of the research process, from evaluating applied environmental to social science questions Daoud and Dubhashi (2023).
However, even under ideal conditions, the predictions of EO-ML models tend to have a smaller variance than the target values, leading to attenuated treatment effects and, thus, potentially misleading conclusions (Berglund, 2012).
Recent research has developed methods that can debias the upstream EO-ML model, but they require either fresh data in the downstream Olofsson et al. (2013); Egami et al. (2023); Lu et al. (2024) or changing the way the model is trained Ratledge et al. (2022). For example, relying on such fresh data, Angelopoulos et al. (2023) proposes Prediction-Powered Inference (PPI), a method that corrects any ML model for downstream inference. Similarly, Egami et al. (2023) achieves the same goal but with better sample efficiency de Pieuchon et al. (2025), but it also requires knowing the probability of observing a sample point. While these methods offer a path to integrate upstream and downstream analyses, their success depends on the sometimes availability of newly collected treatment-predictor-outcome tuples in the downstream causal step or knowledge of the data labeling process (Egami et al., 2023). Often, such data are prohibitively costly or unethical to collect.
Ratledge et al. (2022) addresses the challenge of attenuation bias and the unavailability of data in the downstream by adjusting the model’s loss function instead. During training, they penalize quintile-level discrepancies between observed and predicted means, encouraging predictions to match the true outcome distribution more closely across its support. Although their solutions do not need additional downstream data, they require modifying the training pipeline, which is likely computationally expensive.
Thus, to enable broader applicability of ML models, the upstream team needs to create a data product that is agnostic to downstream use, requires no downstream-fresh data, produces unbiased causal estimates (due to attenuation), and is computationally efficient Daoud and Dubhashi (2023).
We address this gap by proposing Tweedie’s correction, a novel method that reduces prediction bias without requiring fresh labeled data. We evaluate this approach against four others: linear calibration correction, Prediction-Powered Inference (PPI), the loss adjustment approach of Ratledge et al. (2022), and naive uncorrected predictions. The Tweedie’s correction paves the way for the same wealth map to be reused across multiple trials without needing to refit a model for each trial. This also removes the requirement for interaction between the upstream machine-learning team creating the data product and any other domain-knowledge team in the downstream phase. Hence the concept: one map, many trials. We derive analytical proofs and demonstrate the empirical performance of all methods in both simulations and real-world applications, using data from the African Demographic and Health Survey (DHS).
Figure 1: Illustrating the proposed debiasing pipeline for causal inference using machine learning predictions from satellite imagery, without requiring fresh ground-truth data.
1 Problem Setup & Related Work
Statistical predictions can systematically shrink estimates toward the central modes of the training distribution (Ting, 2024a). This phenomenon is rooted in optimal error-minimization estimation (cf. Stein’s paradox), where a model “hedges” predictions in regions of greater uncertainty. When facing noise or limited training data, the optimal strategy, in terms of minimizing residuals, is to pull predictions toward denser regions of the training data. For instance, consider an image that appears with an IWI of 80, but the model has observed this outcome only once, compared to thousands of observations with an IWI of around 70. A perfect prediction would be 80 for this data point, but the model will pull this data point’s prediction towards 70 instead to hedge against the risk of overcommitting to a rare value.
When such biased predictions replace true outcomes in downstream causal analyses, the resulting regression-to-the-mean induces attenuation bias, reducing observed treatment group differences even under randomization (Shu and Yi, 2019). The difference in mean predicted outcomes between treated and control units shrinks both group means toward the global average, thereby attenuating the estimated effect.
Regression-to-the-mean bias is particularly acute in upstream tasks with scarce ground truth (due to finite samples) or in high-dimensional input data (due to the curse of dimensionality), forcing greater reliance on error-prone predictions—something especially pertinent in development economics, political science, and global health contexts —where household data is often expensive or unavailable—to increasingly substitute upstream predictions (Kino et al., 2021; Kakooei et al., 2024; Daoud and Johansson, 2024). Moreover, high out-of-sample $R^{2}$ does not eliminate this bias: models can achieve strong accuracy yet yield attenuated causal estimates (Ting, 2024b).
Existing remedies, such as PPI, which subtracts a rectifier estimated from fresh ground-truth pairs at the causal-inference stage (Angelopoulos et al., 2023), or loss-function adjustments that penalize quintile-level errors (Ratledge et al., 2022). However, these approaches require that the upstream team adapts the data for downstream analysis, slowing or limiting altogether "one map, many trials" efforts.
To address these limitations, we propose a Tweedie-inspired method. This Tweedie method debiases the shrinkage effect in ML-predicted outcomes, enabling downstream researchers to accurately recover causal effects from the upstream data product. Our Tweedie approach treats the ML model as a black box: we assume access only to its out-of-sample predictions and do not rely on retraining or modifying the model. Under this setting, we demonstrate that shrinkage can be reversed, either by learning a global calibration correction from held-out labeled data or by estimating a local correction using ideas from Tweedie’s formula, initially described by Robbins and Tweedie in 1956. Both strategies aim to recover unbiased treatment effect estimates from biased predictions, without requiring fresh ground-truth labels at the time of intervention. We benchmark our Tweedie adaptation to existing methods, which are compared in Table A.III.1 based on their associated assumptions and data/training requirements. As previously mentioned, the only method we are unable to deploy in our setting is the DSL method, as the probability of DHS samples is a function of unknown aid disbursement strategies.
Setup
Our pipeline assumes access to two datasets. First, in the upstream, a labeled dataset with training points (with indices in $\mathcal{I}_{\textrm{Train}}$): $D_{\text{Train}}=\{(\bm{X}_{i},Y_{i}):i=\mathcal{I}_{\textrm{Train}}\},$ where $\bm{X}_{i}\in\mathcal{X}$ denotes covariates (e.g., satellite imagery) and $Y_{i}$ is the true, ground-truth outcome (e.g., IWI poverty level as from the DHS).
Next, in the downstream phase, we assume that different teams have assembled collections of trial datasets indexed by $s$, corresponding to different randomized controlled trials, quasi-experiments, or geographic partitions, with a specific treatment (or action) of interest indexed by treatment setting: $D_{\text{Trial $s$}}=\{(\bm{X}_{i},\widehat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial $s$}},s\in\mathcal{S}\}.$ The data product associated with $Y_{i}$ is general (e.g., all-Africa), but investigators may want to use that data in multiple causal settings, $s$, each with a different treatment, $A_{is}\in\{0,1\}$. The outcome $Y_{i}$ is not directly observed for all $i$. Instead, inference will need to take place through the noisy variable, $\hat{Y}_{i}$. We henceforth remove dependence on $s$ for simplicity.
A key idea henceforth is that the upstream team—the one machine-learning team—creates an agonstic dataset (Grimmer et al., 2022), meaning that this team’s data product should produce unbiased causal estimates for any number of applications (trials) for the variety of teams downstream. This implies that there is no communication between up and downstream teams–hence the firewall between the two phases.
Figure 2: The left is a calibration plot of predicted versus true IWI values at held-out DHS clusters. The orange line shows that the predictions exhibit attenuation bias, resulting in overestimation of wealth for poorer locations and underestimation for wealthier ones. The dashed lines represent the ideal relationship. The right plot shows the residuals.
2 Methodology
In this section, we present a suite of methods designed to debias treatment effect estimates that rely on ML-predicted outcomes. Our overarching goal is to mitigate regression-to-the-mean dynamics in the downstream analysis, improving estimates of causal effects or population-level statistics. We first more formally describe two existing solutions, PPI and the method of Ratledge et al. (2022), before introducing two no-new-data debiasing approaches—Linear Calibration Correction (LCC) and Tweedie’s Correction—that use held-out data (without requiring fresh labels in the causal step) to correct shrinkage bias.
PPI proposes a framework for valid statistical inference when a large set of unlabeled observations is paired with a predictive model’s outputs (Angelopoulos et al., 2023). Concretely, letting $\theta^{*}$ be the target parameter (e.g., a population mean) and $f$ denote the model that maps covariates $\bm{X}_{i}$ to predicted outcomes $f(\bm{X}_{i})$, PPI forms an estimator by adding a rectifier term—an adjustment for the average discrepancy between $Y_{i}$ and $f(\bm{X}_{i})$ observed in a smaller, labeled dataset. An estimate for $\theta^{*}$ is computed using an “imputed” version based on unlabeled data ($\{\bm{X}_{i}\}_{i={n+1}}^{N}$), then corrected by subtracting the bias estimated on the labeled subset. For mean estimates, this adjustment is:
$\hat{\theta}^{\textrm{PPI}}=\frac{1}{N-n}\sum_{i={n+1}}^{N}f(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}\left(f(\bm{X}_{i})-Y_{i}\right)$
Because the rectifier is estimated on ground-truth pairs $\{\bm{X}_{i},Y_{i}\}_{i=1}^{n}$, the procedure maintains valid coverage for $\theta^{*}$ in large samples, provided that the predictive model $f$ is fixed (i.e., not retrained on the labeled data). Unlike our approach, however, PPI crucially relies on gathering labeled observations in the downstream phase. Such labels come from high-quality surveys coordinated by USAID across the African continent and cost millions of US dollars to execute. Given the recent funding cuts to USAID, it is unlikely that new and fresh DHS data points will be collected. Hence, there is an acute need for unbiased data collection, which should be fully handled in the upstream phase. We will nonetheless include PPI correction in our analysis with the goal of showing what would have happened to our estimates had fresh data actually been available downstream.
Another approach to remedy attenuation bias is to adjust the loss function: Ratledge et al. (2022) propose a specialized upstream training procedure that directly penalizes quintile-level discrepancies between observed and predicted means, thereby encouraging $\hat{Y}_{i}$ to match the true distribution of $Y_{i}$ more closely across the support of the outcome distribution. Specifically, this approach modifies the model’s objective function, punishing a quintile having a higher MSE than the others:
$\mathcal{L}=\mathrm{MSE}+\lambda_{b}\hat{E}_{b},$
MSE is the mean squared error, $\lambda_{r}L_{2}$ is the L2 regularization term, and $\lambda_{b}\hat{E}_{b}$ is the quintile-specific bias loss, with $\hat{E}_{b}=\max_{j}(\hat{B}_{j}(\hat{f})^{2})$ and $\hat{B}_{j}(\hat{f})=\mathbb{E}[\hat{f}(\bm{X}_{i})-Y_{i}\mid Y_{i}\in Q_{j}]$. By design, this approach reduces the likelihood that group-level estimates of causal effects are attenuated when $\hat{Y}_{i}$ is used in downstream analyses. In contrast, our methods operate as post-hoc corrections on out-of-sample predictions from a black-box model, requiring no specialized retraining or customized loss function that might reduce overall predictive performance.
2.1 Linear Calibration Correction (LCC)
For LCC, we approximate the relationship between EO-ML predictions and ground-truth wealth on the held-out calibration set by a linear transformation:
$\hat{Y}_{i}=kY_{i}+m,$
where $0<k\leq 1$ reflects shrinkage toward the mean and $m\in\mathbb{R}$ is the intercept, representing the mean fixed bias. This linearity might not always be empirically satisfied, but occurs regularly in our data (see, e.g., Figure A.III.3). When these predictions are used in place of true outcomes to estimate causal effects, this shrinkage attenuates estimated treatment effects.
Let $\tau=\mathbb{E}[Y_{i}\mid A_{i}=1]-\mathbb{E}[Y_{i}\mid A_{i}=0]$ denote the true Average Treatment Effect (ATE). Using $\hat{Y}_{i}$ instead of $Y_{i}$ for ATE estimation yields the bias shown in Proposition 2.1.
Proposition 2.1 (Attenuation bias under linear shrinkage)
Let $Y_{i}$ be the true outcome and $\hat{Y}_{i}$ an ML prediction obeying the linear calibration model:
$\hat{Y}_{i}\;=\;kY_{i}+m+e_{i}\quad(0<k\leq 1,\;m\in\mathbb{R}),$
with $k$ and $m$ constants, and $e_{i}$ mean 0 noise. Let $A_{i}\in\{0,1\}$ denote a binary treatment that is randomly assigned ($A_{i}\perp Y_{i}(a)$ for $a\in\mathcal{A}$). Define the true average treatment effect (ATE) $\tau:=\mathbb{E}[Y_{i}(1)]-\mathbb{E}[Y_{i}(0)]$ and the estimate, $\hat{\tau}$, from difference in predicted expected outcome:
$\hat{\tau}=\mathbb{E}[\hat{Y}_{i}\mid A_{i}=1]-\mathbb{E}[\hat{Y}_{i}\mid A_{i}=0].$
Then $\mathbb{E}[\hat{\tau}]\;=\;k\,\tau.$ That is, the ATE estimated on the predictions is multiplied by the shrinkage factor $k$ and is therefore attenuated toward 0 whenever $k<1$.
Proof sketch.
Take expectations using linearity:
$\hat{\tau}$ $=\mathbb{E}[\hat{Y}_{i}\mid A_{i}=1]-\mathbb{E}[\hat{Y}_{i}\mid A_{i}=0]$
$=\mathbb{E}[kY_{i}+m\mid A_{i}=1]-\mathbb{E}[kY_{i}+m\mid A_{i}=0]=k\tau.$
Replacing observed with potential outcomes follows from randomization. The estimated ATE is biased downward by a factor of $k$.
Now, to adjust for this bias, we will estimate $k$ and $m$ on a held-out calibration set from the training data by regressing $\hat{Y}_{i}$ on $Y_{i}$ (i.e., fitting the linear model $\hat{Y}_{i}=\hat{k}Y_{i}+\hat{m}+e_{i}$, where $e_{i}$ is the residual). We then invert this fitted relationship (solving for $Y_{i}$) to obtain a calibrated estimate, given by $\hat{Y}^{L}_{i}=\frac{\hat{Y}_{i}-\hat{m}}{\hat{k}}$, assuming residuals $e_{i}$ with mean zero conditional on covariates, independent of treatment $A_{i}$. This calibrated estimator restores the scale of the outcome variable, yielding Proposition 2.2:
Proposition 2.2 (Consistency of the LCC-corrected estimator)
Under the conditions of Proposition 2.1, estimate $(k,m)$ on an independent calibration sample and set:
$\hat{Y}^{L}_{i}\;:=\;(\hat{Y}_{i}-\hat{m})/\hat{k},\;\ \hat{\tau}^{L}\;:=\;{\mathbb{E}}[\hat{Y}^{L}_{i}\mid A_{i}=1]-{\mathbb{E}}[\hat{Y}^{L}_{i}\mid A_{i}=0].$
[…]
Proof sketch.
[…]
$\hat{Y}_{i}^{L}$ $=[[kY_{i}+m+e_{i}]-\widehat{m}]/\hat{k}$
$=(k/\hat{k})Y_{i}+(m-\hat{m})/\hat{k}+e_{i}/\hat{k}\to Y_{i}+e_{i}/k$
Hence,
$\hat{\tau}^{L}\to\mathbb{E}[Y_{i}+e_{i}/k\mid A_{i}=1]+\mathbb{E}[Y_{i}+\epsilon_{i}/k\mid A_{i}=0]=\tau,$
since $\mathbb{E}[e_{i}]=0$ and under under noise independence, $e_{i}\perp A_{i}$. A drawback is that we assume a linear model for the shrinkage; estimating $\hat{m}$ and $\hat{k}$ introduces another potential source of error.
2.2 Tweedie’s Correction for Debiasing
To go beyond the linearity assumption in LCC, we now turn to Tweedie’s formula, which allows us to correct shrinkage bias in machine learning predictions by estimating the true outcome as the prediction plus a term involving the derivative of the log-density of the predictions. We follow Efron (2011), who revised the classical results in the context of selection bias as an automatic bias-correction tool in an empirical Bayes framework. We leverage the same core insight here to adjust for regression-to-the-mean bias in satellite-imputed outcomes when analyzing causal effects.
Setup. We here treat the shrinkage mechanism probabilistically:
[…]
$\varepsilon_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\sigma^{2}),\;\;\varepsilon_{i}\;\perp\;(Y_{i},A_{i}).$
The factor $k<1$ pulls extreme $Y_{i}$ values toward the center, lowering $\operatorname*{\mathrm{Var}}(\hat{Y}_{i})$ whenever $\sigma^{2}$ is not too large. Setting $k=1$ recovers the additive-noise model sometimes used in empirical-Bayes texts (Efron, 2012), so (1) nests the classic case.
This noise model enables us to leverage Tweedie’s formula to recover the posterior mean of the latent outcome, accounting for shrinkage without requiring linearity assumptions.
Posterior mean via a scaled Tweedie identity.
Write $W_{i}:=kY_{i}$. Then $\hat{Y}_{i}=W_{i}+\varepsilon_{i}$ with the same $\varepsilon_{i}$ as above, and the Tweedie score identity (Efron 2011) [^1] For any $Z=\theta+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$, $\operatorname*{\mathbb{E}}[\theta\mid Z=z]=z+\sigma^{2}\,\partial_{x}\!\log p_{Z}(z)\,$. gives
$\operatorname*{\mathbb{E}}\!\bigl{[}\,W_{i}\mid\hat{Y}_{i}=\hat{y}\,\bigr{]}\;=\;\hat{y}\;+\;\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}}(\hat{y}),$ (2)
where $p_{\hat{Y}}$ is the marginal density of $\hat{Y}_{i}$; Dividing by $k$ recovers the latent outcome:
${\;\operatorname*{\mathbb{E}}\!\bigl{[}\,Y_{i}\mid\hat{Y}_{i}=\hat{y}\,\bigr{]}\;=\;\frac{\hat{y}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}}(\hat{y})}{k}.\;}$ (3)
For causal estimation, the relevant posterior is conditional on treatment: $\operatorname*{\mathbb{E}}[Y_{i}\mid\hat{Y}_{i}=\hat{y},A_{i}=a]=\frac{\hat{y}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}=a}(\hat{y})}{k}$, since the noise is independent of $A_{i}$ but treatment shifts the distribution of $Y_{i}$.
By expressing the posterior mean of the scaled latent variable in terms of the score function, we obtain a convenient identity that underlies the Tweedie adjustment.
Debiased point estimate.
Equation 3 motivates the Tweedie-adjusted outcome:
$\widetilde{Y}_{i}\;:=\;\frac{\hat{Y}_{i}+\sigma^{2}\,\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}}(\hat{Y}_{i})}{k},$ (4)
which depends solely on the prediction $\hat{Y}_{i}$, treatment $A_{i}$, and two nuisance parameters, $k$ and $\sigma^{2}$.
With this adjustment defined, we now establish its unbiasedness properties under the assumed model.
Proposition 2.3 (Unbiasedness)
Under model (1), $\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}\mid\hat{Y}_{i},A_{i}]=\operatorname*{\mathbb{E}}[Y_{i}\mid\hat{Y}_{i},A_{i}]$ and hence by iterated expectations, $\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}]=\operatorname*{\mathbb{E}}[Y_{i}]$.
Proof Sketch.
Take conditional expectations in (4) and apply the group-specific version of (3) and iterated expectations.
Having established the unbiasedness of the Tweedie-adjusted outcome both conditionally and unconditionally, we now leverage this property to demonstrate that a simple difference-in-means estimator based on these adjusted outcomes yields an unbiased estimate of the average treatment effect under random assignment.
Proposition 2.4 (Tweedie-corrected ATE)
Let $\hat{\tau}^{T}:=\hat{\operatorname*{\mathbb{E}}}[\widetilde{Y}_{i}\mid A_{i}=1]-\hat{\operatorname*{\mathbb{E}}}[\widetilde{Y}_{i}\mid A_{i}=0],$ where $\hat{\operatorname*{\mathbb{E}}}$ denotes the sample mean. Under (1) and random assignment, $\operatorname*{\mathbb{E}}[\hat{\tau}^{T}]=\tau$.
Proof Sketch.
$\operatorname*{\mathbb{E}}[\widetilde{Y}_{i}\mid A_{i}=a]=\operatorname*{\mathbb{E}}[Y_{i}\mid A_{i}=a]$ by Prop. 2.3; subtract.
Figure 3: Illustration of Tweedie’s correction mechanism, showing how the posterior mean adjusts for shrinkage bias in predicted outcomes under Gaussian noise.
Estimation
The parameters $k$ and $\sigma^{2}$ come from a held-out calibration set $\mathcal{I}_{\text{cal}}\subset\mathcal{I}_{\text{Train}}$ upstream; scores are estimated downstream:
  1. 1.
    Shrinkage factor $k$. Regress $\hat{Y}_{i}$ on $Y_{i}$: $\hat{Y}_{i}=\hat{k}\,Y_{i}+\hat{m}+\text{resid}$. We keep the slope $\hat{k}$ (discarding the intercept, absorbed into $Y_{i}$).
  2. 2.
    Noise variance $\sigma^{2}$. Use the residuals: $\hat{\sigma}^{2}=|\mathcal{I}_{\text{cal}}|^{-1}\sum_{i\in\mathcal{I}_{\text{cal}}}(\hat{Y}_{i}-\hat{k}Y_{i})^{2}.$
  3. 3.
    Group-specific score functions $\partial_{\hat{y}}\log p_{\hat{Y}\mid A_{i}=a}(\hat{y})$. Downstream, for each $a\in\{0,1\}$, fit a smooth density $\widehat{p}_{\hat{Y}\mid A_{i}=a}$ to $\{\hat{Y}_{i}:i\in\mathcal{I}_{\text{Trial}},A_{i}=a\}$. Trim or regularize the derivative where $\widehat{p}_{\hat{Y}\mid A_{i}=a}$ is small.
The operational formula passed downstream is in the end:
$\widehat{\widetilde{Y}}_{i}=\frac{\hat{Y}_{i}+\hat{\sigma}^{2}\,\partial_{\hat{y}}\log\widehat{p}_{\hat{Y}\mid A_{i}}(\hat{Y}_{i})}{\hat{k}}.$
Relation to LCC.
When the score term tends to zero (e.g., occurring when the variance on outcomes, $Y_{i}$, grows large), $\widetilde{Y}_{i}=\hat{Y}_{i}/k$—exactly an LCC correction. Tweedie’s adjustment, therefore, refines LCC with a local, data-driven addend that vanishes in regions where $p_{\hat{Y}\mid A_{i}}$ is flat and grows where shrinkage is most severe. For simplicity, we will, for all empirical Tweedie analyses henceforth, set $k=1$ to isolate the performance contribution of the local, data-driven adjustment (versus that from the LCC-type correction).
3 Experiments with Simulated Data
Simulation Design.
Our simulation study benchmarks all debiasing methods and monitors the upstream and downstream analyses while maintaining the firewall between them. The simulation uses a canonical directed acyclic graph (DAG), shown in Figure 4. It is canonical in the sense that the EO-ML literature tends to assume this structure (Sakamoto et al., 2025). It illustrates the various research phases and their corresponding data availability.
In phase 0, the data-generation process (approximating reality) unfolds in the direction of the arrows. In the upstream phase, only the blue-nodes data are available, with finite samples of $Y_{i}$, to produce $\hat{Y_{i}}$ for the entire population $N$. In the downstream, only red nodes are available: $\hat{Y_{i}}$ is now available for $N$, jointly with $C_{i}$ and $A_{is}$. They enable inference on the estimand $\tau_{s}$. (We suppress the index $s$ henceforth.) Appendix IV delineates our simulation further.
Figure 4: A panel illustrating data availability in the two phases.
0. Reality (data generation) shows the complete causal DAG.
1. Upstream phase: $Y_{i}$, $\mathbf{X}_{i}$, and $\hat{Y}_{i}$ are observed (blue).
2. Downstream phase: $A_{is}$, $C_{i}$, and $\hat{Y}_{i}$ are available (red.)
Simulation Results.
We focus on two key objectives: minimizing error in the upstream task and minimizing causal bias in the downstream task. To assess error, we measure how closely the predicted values match the true values of $\tau$, using Mean Absolute Error (MAE). A lower MAE indicates a more accurate estimation of the outcomes, and thus, a more accurate downstream ATE estimate. To assess bias, we examine the extent to which the upstream outcome prediction errors vary systematically when changing the magnitude of $\tau$. This examination entails regressing the predicted $\hat{\tau}$ values on the true $\tau$ values and analyzing the calibration slope. If the estimates are unbiased, the slope should be close to one. Deviations from one indicate systematic over-/underestimation across the range of $\tau$.
We can formally test deviations from 111 with an $F$-test: $F$-statistics from the above regression quantify the strength of the relationship between the predicted treatment effects and the true treatment effects. These $F$-statistic results (Table 1) indicate that Tweedie’s correction is the only method that yields ATE estimates for which we cannot reject the null hypothesis that these estimates are unbiased. Figure A.III.1 visualizes these slope results.
Method MAE Slope $F$-statistic $p$-value
Tweedie’s 0.04 0.995 1.22 0.303
LCC 0.05 1.008 5.75 0.006
PPI (10%) 0.19 0.985 6.441 0.003
Ratledge 0.37 0.641 32.93 0.000
Naive 0.48 0.535 6420.43 0.000
Table 1: Simulation performance metrics. $F$-statistics and $p$-values are the statistical test for evaluating the null hypothesis that the slope value is 1. Failure to reject indicates that the slope estimate intervals contain the true ATE.
4 Experiments with Real Data
In this section, we conduct two types of downstream-real experiments. First, we will evaluate how well our debiasing methods succeed in descriptive tasks, especially those involving group aggregation. Second, we will apply our procedure to causal tasks. These two tasks cover the main use cases of our EO-ML-driven poverty data.
The full dataset consists of $n\approx$ 69,000 geo-located survey clusters from the Demographic and Health Surveys (DHS) across 30 African countries (2009–2020), where each cluster comprises 20–30 households sampled via stratified two-stage design (primary sampling units selected with probability proportional to population size, followed by random household selection). We partitioned the data into five random folds for cross-validation (CV). For each run, four folds were used as upstream data (three for training, one for calibration): $D_{\text{Train}}=\{(\mathbf{X}_{i},Y_{i}):i\in\mathcal{I}_{\text{Train}}\}$. Downstream trial datasets $D_{\text{Trial }s}$ were drawn from the remaining, fifth fold and geo-referenced aid interventions, $s$, ensuring no overlap with training data.
Here, $Y_{i}$ denotes the International Wealth Index (IWI), a continuous asset-based wealth score (0–100) aggregated at the cluster level to mitigate noise (Burke et al., 2021; Zhu et al., 2025). The covariate array $\mathbf{X}_{i}$ consists of Landsat satellite imagery (30m resolution, six multispectral bands) centered on each DHS cluster (6.72 km × 6.72 km tiles), processed by taking the three-year median over cloud-free pixels leading up to the survey date.
Two models were fitted for each fold, one using MSE and one using the loss introduced by Ratledge et al. (2022). The models shared the same ResNet-50 architecture and were initialized with weights pretrained on the SSL4EO-L dataset (Stewart et al., 2023). For more details about model training, see Appendix II. The resulting MSE models all have an $R^{2}$ score between 0.74 and 0.76 on their hold-out fold, and an average slope of 0.79 . Their combined predictions can be seen in Figure 2. The Ratledge models have a slope of 0.99, but at the cost of reduced $R^{2}$ scores between 0.67 and 0.70.
Estimating Region Means with Bias Adjustment.
In our descriptive task, we estimate conditional means, averaging wealth by local second-level administrative region (ADM2), roughly corresponding to municipalities or counties. We group the clusters in the held-out downstream fold by ADM2 region and compare the mean observed IWI with the predicted IWI with and without the different methods of bias correction. These quantities are subject to the same shrinkage bias as causal estimates, and our framework applies equally well in this setting, enabling us to assess how effectively we correct biased group-level estimates.
Descriptive Results.
As expected, the MSE of the naive model exhibits shrinkage behavior: it tends to overpredict the wealth of poorer regions and underpredict the wealth of richer regions. As shown in Figure 5, Tweedie’s correction accurately compensates for this, not only resulting in less biased estimates (regression slope of 0.83 to 0.90), but also decreases the MAE from 2.67 to 2.39. In fact, if we consider only regions with at least 100 clusters, the Tweedie-corrected estimates exhibit virtually no shrinkage bias, with a slope of 0.99. Appendix III delineates this analysis.
Figure 5: Mean residuals and Tweedie’s corrections for each ADM2 region in the dataset. regions that are overestimated in the left plot (blue) are shown to be correctly adjusted down in the right plot (red), and vice versa.
Empirical Analysis of Aid Interventions.
To assess the practical utility of our correction methods in a setting where the true treatment effect is unknown, we construct an evaluation framework based on real-world development projects. Specifically, we leverage the same dataset of geo-referenced interventions funded by the World Bank and the Chinese government as was used in Malik et al. (2021); Conlin (2024). Each project is tagged with an associated aid sector $s\in\mathcal{S}$ (e.g., Health, Water Supply and Sanitation, or Women in Development), which defines a downstream trial $D_{\text{Trial }s}=\{(\bm{X}_{i},\widehat{Y}_{i},A_{is}):i\in\mathcal{I}_{\textrm{Trial }s}\}$. Here, $A_{is}\in\{0,1\}$ indicates exposure to a $s$ intervention at location $i$, $\bm{X}_{i}$ are satellite features, and $\widehat{Y}_{i}$ are IWI predictions from the upstream-trained model. Figure A.III.2 depicts the spatial distribution of treatments across the continent.
Since the treatment effects of these interventions are unknown, we deem this analysis a test of external validity by comparing EO-ML predicted wealth values with observed ones. Our goal is to assess whether our bias-corrected predictions can reflect causal estimates similar to those from observed data.
To avoid label leakage and emulate real-world constraints, we perform a two-stage evaluation. We begin by randomly partitioning the available DHS survey data into two disjoint sets: an upstream set for training and calibrating the EO-ML model, and a downstream set used solely for evaluation. The model is trained on the upstream data, without any exposure to the intervention data or knowledge of future evaluation criteria. Shrinkage corrections (e.g., LCC or Tweedie’s correction) are also estimated on this upstream set.
We match project locations to the corresponding ADM2-level administrative regions, roughly equivalent to districts or municipalities. For a given funder-sector pair (e.g., Health projects funded by the World Bank), we identify all ADM2 regions containing at least one relevant intervention site. We then consider villages surveyed in the downstream DHS data within those regions, 3 to 8 years after the intervention, as the treated group. Villages in the same surveys, but outside the treated ADM2 regions, serve as controls.
For each funder-sector combination, we estimate the average treatment effect as the difference in mean outcomes between treated/control villages. We compute this using both observed IWI values and the model-predicted (and correction-adjusted) IWI values.
Causal Results.
See Figure 6 for results. Each point is a causal study. While the $x$-axis shows what would have happened when using the label $Y_{i}$ in these studies, the $y$-axis shows the corresponding causal effect using $\hat{Y_{i}}$ with its corresponding adjustment approach. Orange lines check the correlation among these data. A score of one is perfect. The analysis shows that Tweedie’s correction (panel d) comes close to this desiderata, with a correlation of 0.9980.9980.998; and it does so with the lowest MAE of 0.570.570.57. Had we used the naive predictions (panel a), the correlation would be 0.9580.9580.958 and the MAE would be 0.690.690.69. Assuming the PPI set up, with 10% fresh data available, we would have achieved a 0.9810.9810.981; and it does so with the lowest MAE of 0.730.730.73. The PPI performance is, however, a function of available data in the downstream, and panel f shows that with increasing fresh data, the PPI correction will outperform our Tweedie. But PPI is fresh-data hungry, and thus, this approach would come at a cost of deploying more costly surveys in Africa—costing downstream teams millions of USD de Pieuchon et al. (2025).
Figure 6: (a–e) Estimated treatment effects of aid interventions using unadjusted predictions, which display attenuation bias due to shrinkage. (f) PPI outperforms Tweedie’s correction in MAE bias reduction on larger hold-out sets, but requires many ground-truth labels.
5 Discussion and Conclusion
Our experiments show that post-hoc, data–free corrections can eliminate the lion’s share of the shrinkage-induced attenuation that plagues causal analyses based on machine–learning predictions. In controlled simulations where the true ATE is known, Tweedie’s correction reduces mean–absolute error by an order of magnitude and, crucially, restores calibration: the fitted slope of $\hat{\tau}$ on $\tau$ is indistinguishable from one (Table 1, Figure A.III.1). The linear calibration correction also performs nearly as well despite its simplicity, suggesting that much of the bias arises from a first-order scaling distortion that can be identified on a small held-out calibration set. When applied to real DHS imagery, both corrections markedly narrow the gap between predicted and observed country-level means (Figure A.III.3), and in a quasi-experimental evaluation of large-scale aid projects, they bring estimated ATE into line with those obtained from ground-truth survey data (Figure 6).
For future research, we see three immediate research paths. First, formalize a theory of adaptive corrections that tunes the strength of Tweedie’s adjustment to local density estimates, thereby taming heteroskedasticity. Second, extend the methodology to multi-task settings where outcomes such as health, education, and nutrition are predicted jointly—an arena. Third, integrate causal discovery into the pipeline, allowing the system to learn when a variable is about to “debias” and find post-treatment leakage Daoud et al. (2022). That would prevent the analyst from inadvertently resurrecting bias. It is important to note, however, that our methods specifically address attenuation bias arising from shrinkage and do not resolve all forms of bias in machine learning predictions; therefore, users should not regard the corrected predictions as fully unbiased. $\square$